### v1

# revtex4-2Repair the float

Abstract
Motivated by the recent realization of space-borne Bose-Einstein Condensate (BEC) under micro-gravity conditions, we extend the understanding of ultracold dipolar bosonic gases by exploring their behavior in a novel trapping configuration known as the “bubble trap” topology. Utilizing the three-dimensional numerical simulations within the extended Gross-Pitaevskii framework, we unveil diverse ground state phases in such a static curved topology. Subsequently, we investigate the influence of rotation on a dipolar BEC confined to the surface of a spherical bubble. Our findings reveal that the rotation of a bubble trap with certain rotation frequencies can modify the effective local dipole-dipole interaction strength, leading to the induction of supersolidity and the formation of quantum droplets. In addition, we demonstrate that a bubble trap can sustain high circulation and the flow also persists for a longer time. Significantly, adjusting the rf detuning parameter allows the condensate to achieve hypersonic velocity. Finally, we explore the impact of drastic change in the topological nature of the trap on the rotating dipolar BEC, transitioning from a filled shell trap to a bubble trap and vice versa.
Introduction.- The exploration of trapped atomic (Bose-Einstein Condensate) BEC has traditionally been steered by the investigations of atom-atom interaction strength, dimensionality, and topological nature of the trapped geometry. The ability to tune the interaction strength between the atoms of ultracold atomic gases [1, 2, 3, 4] gives rise to a plethora of fascinating physical phenomena, including the superfluidity [5], formation of quantized vortices [6, 7, 8, 9, 10], solitons [11, 12, 13, 14, 15], the superfluid to Mott insulator phase transition in an optical lattice [16, 17] and more recently the formation of a quantum droplet [18, 19, 20, 21] and supersolid phase [22, 23, 24]. The signature of supersolidity was initially detected in two experiments utilizing BECs within an optical cavity [25] and with mixtures featuring spin-orbit coupling [26]. In recent years, the realization of quantum droplets and supersolid states in dipolar bosonic gases of rare-earth Dy [27, 21, 20] and Er [28, 29, 22] atoms has attracted significant attention. Theoretical and experimental studies over the past few years reveal the appearance of roton excitation [30, 31, 29, 32, 33, 34, 35, 36, 37], formation of the self-bound quantum droplets [20, 19, 38, 39] and supersolid state in quasi one [40, 41, 24, 42, 43] and two-dimensional [33, 44, 45, 46, 47, 48] harmonically trapping confinements.
The trap geometry of a BEC also plays a crucial role in shaping various intriguing physical phenomena. Examining BECs in one dimension (1D) provides a deeper understanding of fermionization [49] and many-body systems in non-equilibrium conditions [50]. In two dimensions (2D), the study of BECs has shed light on quasi-condensation, the Berezinskii-Kosterlitz-Thouless (BKT) transition [51, 52] and the formation of a two-dimensional vortex lattice [53]. Condensates in double-well and optical lattice potential have numerous applications, such as matter-wave interferometry [54, 55] and spin squeezing [56]. A rotating toroidal or ring trap can maintain a continuous superfluid flow with substantial angular momentum which also persists for a longer time [57, 58, 59, 60, 61].
Recent advancements in the precision and sophistication of controlling the trapping mechanisms for ultracold atomic gases led to the formation of a novel trapping geometry resembling that of a bubble, where all the atoms in the condensate are confined to the surface of the bubble [62, 63, 64]. The unique topology and curved geometry of this bubble trap open a promising new avenue for intriguing research endeavors, including the investigations into the influence of curvature on various topological defects and their dynamics [65, 66, 67, 68], exploration of new collective modes [69, 70, 71], examination of matter-wave interference [69], comprehension of the persistent flow of the superfluid [72], the emergence of supersolid phase in bubble trap [73], and the study of dimensional crossover from harmonic to thin bubble trap configurations [70, 74, 75]. A recent study reveals the formation of equatorial waves in a rotating bubble-trapped BEC [76].
Motivated by this optimistic trapping configuration and the emergence of quantum droplets and supersolid phases in dipolar BEC, in this Letter, we study the effect of dipolar BEC confined on the surface of a bubble trap and the influence of the trap rotation on it. Remarkably, we find that the rotation of the bubble trap with certain rotation frequencies modifies local relative dipole-dipole interaction (DDI) strength, which could induce supersolidity and nucleation of quantum droplets in an otherwise superfluid state. Additionally, we note the ability to sustain a state with exceptionally high circulation within this bubble-shaped topology. This flow with a large angular momentum per particle also persists for a longer duration. Moreover, we observe that a rotating bubble trap, characterized by a large radius and strong transverse confinement, leads to the attainment of hypersonic velocity of the condensate. Furthermore, we have also investigated the impact of altering the topology of the trap geometry, transitioning from a filled spherical shell to a thin shell and vice versa on the dipolar BEC.
Formalism.- At zero kelvin temperature, the dipolar BEC of atoms with mass $m$ and magnetic dipole moment $\mu$ is described by the macroscopic wave function $\psi(\mathbf{r},t)$, whose temporal evolution is governed by extended Gross-Pitaevskii equation (eGPE) $i\hbar\dot{\psi}(\mathbf{r},t)=\hat{H}\psi(\mathbf{r},t)$ [77, 28, 78, 79] with the Hamiltonian,
$\hat{H}=\Big{[}-\frac{\hbar^{2}}{2m}\nabla^{2}+V(\mathbf{r})+g|\psi(\mathbf{r},t)|^{2}$
$+\int U_{\rm dd}(\mathbf{r}-\mathbf{r^{\prime}})|\psi(\mathbf{r^{\prime}},t)|^{2}d\mathbf{r^{\prime}}+\gamma(\epsilon_{\rm dd})|\psi(\mathbf{r},t)|^{3}\Big{]}.$ (1)
[…]
$V(\mathbf{r})=m\omega^{2}\sqrt{(\mathbf{r}^{2}-\Delta)^{2}/4+\Theta^{2}}-m\omega^{2}\Theta.$ (2)
Here $\Delta$ represents the detuning between the applied rf field and the participated energy states of the condensate, and $\Theta$ denotes the Rabi frequency between these states. These parameters control the radius and width of the bubble trap. For $\Delta=\Theta=0$ the bubble trap (Eq. (2)) reduces to a well-known spherical harmonic trap. For a fixed value of $\Theta$ and with the larger value of $\Delta(\Delta>\Theta)$ we get a thinner spherical shell potential with a larger radius and it can be approximated as a radially shifted harmonic trap (SHT) $V_{\rm SHT}(\mathbf{r})=\frac{m\omega^{2}\Delta}{2\Theta}(\mathbf{r}-\sqrt{\Delta})^{2}$ [70], as shown in Fig. 1.
Figure 1: Comparison between bubble trap potentials as defined by Eq. (2) with $\Theta=1~{}\mu\rm{m}^{2}$ for different values of $\Delta$. The dashed line plots of respective color demonstrate their closely approximated radially shifted harmonic trap (SHT) potential with the same $\Delta$ values.
Non-rotating ground state.- To begin our exploration of the rotating state in a bubble trap, it is imperative to first delve into the characteristics of the non-rotating states. We examine the ground state configurations by numerically evolving the eGPE in the imaginary time using the split-step Crank-Nicolson method [83]. For definiteness, we consider $N=6\times 10^{4}$ number of ${}^{164}$y atoms confined in a bubble trap with trap frequency $\omega=2\pi\times 100$ Hz in all three directions, and the other parameters $\Delta=36l_{\rm{osc}}^{2}=21.9~{}\mu\rm{m}^{2}$ and $\Theta=0.0441l_{\rm{\rm{osc}}}^{2}$, where $l_{\rm{osc}}=\sqrt{\hbar/m\omega}=0.78~{}\mu\rm{m}$ is the characteristic length scale set by the trap. The ${}^{164}$y atom has the magnetic dipole moment $\mu=9.93\mu_{\rm B}$ and the corresponding dipolar length $a_{\rm{dd}}=131a_{\rm B}$, where $\mu_{\rm B}$ and $a_{\rm B}$ is the Bohr magneton and Bohr radius, respectively.
The different ground state configurations are emerging based on the value of $\epsilon_{\rm{dd}}$, which can be scaled experimentally by tuning the s-wave scattering length through the Feshbach resonance technique [84]. The various ground states are distinguished by their superfluid fraction defined through the non-classical rotational inertia according to $f_{s}=1-I/I_{\rm{cl}}$ [85, 86], where $I$ is the moment of inertia of the system and $I_{\rm{cl}}$ is its classical rigid body moment of inertia. We obtain that the associated solutions for $\epsilon_{\rm dd}<1.35$ are fully superfluid BECs with unmodulated density distribution [see Fig. 2(a)] and a superfluid fraction close to unity ($f_{s}>0.9)$ [see Fig. 2(d) ]. As $\epsilon_{\rm dd}$ is increased to favor the DDI strength over contact interaction, beyond a critical value ($\epsilon_{\rm dd}\approx 1.35$) a spontaneous density modulation is developed due to the softening of the roton mode [31]. In the interval $1.431.35<\epsilon_{\rm dd}<$ 1.43, the ground state displays a modulated density profile [see Fig. 2(b)] and it appears to be a regularly arranged spheroidal-like droplet, elongated along the polarisation direction and immersed in a background of low-density condensate with the superfluid fraction less than unity ($0.90<f_{s}<$ 0.9) [see Fig. 2(d) ]. Further increase in the value of $\epsilon_{\rm dd}$, the overlap between the droplets vanishes, and the ground state exhibits incoherent crystal of isolated droplets [see Fig. 2(c)] with $f_{s}=0$ [see Fig. 2(d) ].
Figure 2: Non-rotating ground state density profiles in the $z=0$ plane (a)-(c) for different values of $\epsilon_{\rm dd}$ representing (a) superfluid ( ), (b) supersolid ( ) and (c) incoherent crystal of isolated droplet ( ) phase in a bubble trap with $\Delta=21.9~{}\mu\rm{m}^{2}$ and $\Theta=0.0441l_{\rm{\rm{osc}}}^{2}$. (d) Shows the superfluid fraction $f_{s}$ as a function of dimensionless parameter $\epsilon_{\rm{dd}}$. The markers in (a)-(d) and the background hue in (d) signify the distinct phases.
Depending on the value of $\epsilon_{\rm dd}$, the ground state in a bubble trap exhibits similar phases as dipolar BEC in a harmonic trap [22, 46], albeit with distinct density distributions. The difference in the density distribution results from the bubble trap confinement, which leads to an exceptional arrangement of dipoles within the system. The dipoles are aligned in a head-to-tail fashion along the polar angle $\theta$ and side-by-side along the azimuthal angle $\phi$. This specific arrangement yields distinct interactions for dipoles positioned along the equator and those at the pole. Equatorial dipoles experience a combination of attractive and repulsive forces, originating from their neighboring dipoles along the equator and those positioned above and below them along the meridian lines, respectively. Conversely, dipoles located at the poles encounter solely repulsive forces from their surrounding dipoles as a result of side-by-side arrangement. Consequently, the droplets adopt a curved shape along the surface of the bubble, exhibiting a spheroidal-like distribution elongated along the polarization direction (See Fig. S1 in Supplementary Material), and the density at the pole remains lower to minimize the overall energy of the system.
Rotating ground state.- Once the ground state phases are established, we investigate the impact of rotation on the ground state configurations at different rotation frequencies within the superfluid phase regime (i.e., $\epsilon_{\rm dd}<1.35$). We determine the rotating ground states by incorporating the angular momentum constraint $-\Omega\hat{L}_{z}$ into the Hamiltonian (Eq. (1)) and evaluating the modified eGPE in imaginary time. Here, $\hat{L}_{z}$ represents the $z$ component of the angular momentum operator, and $\Omega$ denotes the rotation frequency.
Figure 3: Rotation-induced spontaneous density modulation and formation of supersolid-like density distribution at specific ranges of rotation frequencies. (a)-(f) show the rotating ground state density profiles for $\Omega=0.12\omega,0.19\omega,0.33\omega,0.53\omega,0.64\omega,$ and $0.80\omega$ in the $z=0$ plane, respectively. See Fig. S2 in the Supplementary Material for three-dimensional isosurface plots. Other parameters are same as of Fig. 2(a).
For instance, at $\epsilon_{\rm dd}=1.31$ the non-rotating ground state exhibits a superfluid state with a smooth, unmodulated density distribution on the surface of the bubble as illustrated in Fig. 2(a). However, as we introduce rotation at certain rotation frequencies the condensate spontaneously develops density modulation and the system manifests a supersolid-like density configuration with periodic density humps on the curved surface of the bubble interconnected by lower density region [see Fig. 3]. At these rotation frequencies, the DDI energy ($E_{\rm{DDI}}$) not only remains attractive in nature but also exhibits sudden dips in its value, as shown in Fig. 4(a). These dips arise from the modification of the effective local DDI by rotation at specific ranges of $\Omega$ values, emphasizing the pronounced role of DDI. As a consequence of this, the superfluid fraction of the condensate also reduces at these specific ranges of $\Omega$ values, indicating the emergence of a supersolid state with $0.90<f_{s}<$ 0.9 [see Fig. 4(a)]. A larger dip in the value of $E_{\rm{DDI}}$ corresponds to a more significant modulation in density. However, this distinctive feature is not observed in either a rotating dipolar BEC in a harmonic trap or a non-dipolar BEC in a rotating bubble trap. The exclusive arrangement of dipoles due to bubble trap confinement, coupled with anisotropic DDI, makes the rotation-induced supersolid transition energetically preferable compared to the conventional formation of quantized vortices for these specific ranges of $\Omega$ values.
Figure 4: (a) Variation of DDI energy $E_{\rm{DDI}}/N\hbar\omega$ (black line with triangular markers) and superfluid fraction $f_{s}$ (purple line with circular markers) for rotating ground states with trap rotation frequency $\Omega$. The dips at specific ranges of $\Omega$ values resemble the increase in the attractive DDI which leads to the rotation-induced supersolid transition. (b) Shows angular momentum $\langle\hat{L}_{z}\rangle/N\hbar$ of the rotating ground state as a function of the rotation frequency $\Omega$ of the trap for $\epsilon_{\rm dd}=1.31$. The step-like growth (the red solid line is drawn for clarity) signifies the quantization of angular momentum corresponding to the superfluid phase, while the intermediate linear growth indicates a solid body response and characterizes the supersolid state. The insets (c) and (d) show the density distribution in the $z=0$ plane for $\epsilon_{\rm dd}=1.31$ at $\Omega=0.30\omega$ and $\Omega=0.53\omega$, respectively. The light green and the light red background specify the regions of the superfluid and supersolid state, respectively. Other parameters are same as of Fig. 2(a).
The angular momentum of the system is also changed distinctively. In Fig. 4(b), we depict the variation of the expectation value of angular momentum ($\hat{L}_{z}$) per particle of the rotating ground states as a function of the rotation frequency ($\Omega$). Generally, when subjected to rotation, the angular momentum of a superfluid state displays its characteristic step-function behavior with $\Omega$, due to the quantization of circulation [9, 10, 87, 88]. However, in a bubble trap, alongside this step function, a linear increase in angular momentum occurs for specific ranges of $\Omega$ values. This linear growth signifies a solid-body response to the rotation, attributed to the emergence of a supersolid structure.
Remarkably, we also found that the transfer of angular momentum is more efficient in bubble trap confinement than in the harmonic trap, leading to higher angular momentum per particle at lower $\Omega$ values, with vortices nestled inside zero-density regions. This leads to the circulation of the condensate with a substantial angular momentum around the rotation axis. Notably, this flow persists even after halting the rotation of the trap as shown in Fig. S3 in Supplementary Material.
Figure 5: (a) Angular momentum $\langle\hat{L}_{z}\rangle/N\hbar$ of the ground state in the rotating frame shows linear variation with $\Delta$ for $\epsilon_{\rm dd}=1.31,\Omega=0.9\omega$ and $\Theta=0.0441l_{\rm{osc}}^{2}$. With these parameters, the inset (b) show the density isosurfaces of a supersolid state at $\Delta=60.84~{}\mu\rm{m}^{2}$, and the inset (c) shows the density isosurfaces of a droplet state formed by macrodroplets at $\Delta=243.4~{}\mu\rm{m}^{2}$. Note that $\langle\hat{L}_{z}\rangle/N$ is approximate to $14\hbar$ for $\Delta=0$ and reaches close to $276\hbar$ at $\Delta=243.4~{}\mu\rm{m^{2}}$.
Moreover, there is a growing interest in preparing strongly correlated ultracold quantum bosonic gases within the quantum Hall regime [89, 90, 91], but achieving the necessary angular momentum poses a challenge. The rotation frequency required to generate such substantial angular momentum in a harmonically trapped BEC surpasses the deconfinement frequency. However, the ability to generate larger angular momentum in the bubble trap than in the harmonic trap makes it a promising candidate for obtaining a quantum Hall state. Furthermore, in experiments, the $\Delta$ parameter can be tuned by modulating the rf detuning. Increasing the value of $\Delta$ for a fixed rotation frequency $\Omega$ results in a larger moment of inertia of the condensate and facilitates access to larger angular momentum. Fig. 5(a), illustrates the growth of angular momentum with $\Delta$ for a constant rotation frequency $\Omega=0.9\omega$. In our numerical simulations, we vary the value of $\Delta$ from 00 (spherical harmonic trap) to $400l_{\rm{osc}}^{2}=243.4~{}\mu\rm{m^{2}}$ (bubble trap). For $\Delta=243.4~{}\mu\rm{m^{2}}$ and $\Omega=0.9\omega$, we observe that the angular momentum per particle reaches $276\hbar$. Further, due to the attractive DDI, the condensate’s chemical potential decreases, resulting in a decrease in sound velocity $v_{s}=\sqrt{\mu/m}$. Consequently, in this scenario, the local fluid velocity $v=\sqrt{\Delta_{\rm{max}}}\Omega\approx 8.817~{}\rm{mm/s}$ significantly surpasses the sound velocity $v_{s}\approx 0.124~{}\rm{mm/s}$, indicating that the condensate rotates at a hypersonic speed with a Mach number $v/v_{s}\approx 71$. Interestingly, as we increase $\Delta$, the transverse trapping frequency also rises (see Eq. (2) and Fig. 1), causing the rotation-induced droplets to merge with nearby ones and form macrodroplets immersed within a background superfluid [Fig. 5(a1)]. Further increasing the value of $\Delta$ results in the suppression of the density of the background superfluid, leading to the formation of an isolated macrodroplet crystal state [Fig. 5(a2)].
Topological transition.- So far we have reported several intriguing phenomena that can emerge as a result of the unique bubble-shaped topology. In order to delve deeper into the role of the topology of the trap, we thoroughly investigate the impact of dynamical transitioning from a bubble-trapped BEC to a spherical-filled BEC and vice versa by tuning experimentally controlled parameter $\Delta$ with fixed $\Theta=0.0441l_{\rm{osc}}^{2}$.
To initiate this study, we first consider a rotation-induced supersolid state in a bubble trap with $\Omega=0.53\omega$ and $\Delta=21.9~{}\mu\rm{m}^{2}$. Starting from this ground state configuration, we linearly decrease [^1] Irrespective of the initial state and trap rotation frequency, a sudden transition from a bubble trap to spherical harmonic trap results in the emergence of matter-wave interference and reduction in the condensate’s lifetime. See Fig. S4 in the Supplementary Material. the parameter $\Delta$ to a final value of $\Delta=0$, representing a spherical harmonic trap. This modulation of $\Delta$ is executed over a time period of $50~{}\rm{ms}$ and after which we let the condensate evolve for a certain time, keeping $\Omega$ constant. Initially, owing to the bubble-shaped trap geometry and influenced by rotation ($\Omega=0.53\omega$), we observe that the condensate exhibits a supersolid state [Fig. 6(a1)] and is spinning about the rotation axis of the bubble trap. This results in a relatively large angular momentum ($\langle\hat{L}_{z}\rangle/N\approx 18\hbar$) compared to the initial ground state [Fig. 6(b1)] in a harmonic trap ($\langle\hat{L}_{z}\rangle/N\approx 1.54\hbar$). During the linear decrease in the value of $\Delta$, the rotational symmetry ensures $[\hat{H},\hat{L}_{z}]=0$, implying that the expectation value of $\hat{L}_{z}$ remains conserved. Consequently, the angular momentum of the condensate, stemming from its rotation about the rotation axis in the bubble trap, is transferred to the condensate within the spherical harmonic trap in the form of topological excitations of vortices. Due to the significant angular momentum, the condensate’s radius expands and accommodates a large number of vortices distributed in a spiral pattern [see Figs. 6(a2) and 6(a3)].
As we stated, achieving a strongly correlated quantum Hall regime necessitates a large angular momentum, surpassing the deconfinement frequency limit for harmonically trapped ultracold gases. However, instead of following a direct dynamic stirring path, one can follow the abovementioned protocol. Starting from a rotating bubble trap with large $\Omega$ and $\Delta$, dynamically reduce the parameter $\Delta$ to zero. Given the ability to attain substantial angular momentum within the bubble trap and the conservation of angular momentum, this approach has the potential to facilitate the realization of the desired quantum Hall regime for a harmonically trapped BEC.
[…]
We carry out the same dynamics in the opposite direction by changing the topology from a filled sphere ($\Delta=0$) to a hollow shell ($\Delta=21.9~{}\mu\rm{m}^{2}$). Initially, with the specified parameter value of $\epsilon_{\rm{dd}}=1.31$, the condensate within the harmonic trap exhibits a superfluid state. Under the influence of trap rotation ($\Omega=0.53\omega$), the condensate demonstrates four off-center vortices at $t=0~{}\rm{ms}$ [see Fig. 6(b1)], a characteristic feature of the superfluid phase. As we increase the value of $\Delta$, the density at the center diminishes and the vortices become nested at the center within a zero-density region. As a consequence of the conservation of angular momentum, the angular momentum of the vortices transforms into the spinning motion of the condensate about the rotation axis. Furthermore, analogous to the ground state of a bubble trap rotating with the same frequency ($\Omega=0.53\omega$) [see Fig. 3(d)], the increase in local DDI induces density modulation in the condensate, leading to the formation of a supersolid state [see Figs. 6(b2) and 6(b3)].
Conclusions.- We have systematically examined the ground state phases of a dipolar BEC by varying the parameter $\epsilon_{\rm dd}$ within a bubble trap. Our investigation reveals that rotation applied to the bubble trap with certain ranges of rotation frequencies induces a transition from superfluid to supersolid phase, achieved through the modification of local DDI strength. Additionally, we have observed that condensate confined in a bubble trap can sustain higher circulation for an extended duration. Notably, our findings also demonstrate the emergence of macrodroplets rotating with hypersonic speed at higher rotation frequencies and larger values of $\Delta$. Furthermore, our analysis of the hollowing transition, a transition from a filled spherical BEC to a hollow BEC in the rotating frame, confirms that the condensate undergoes modulation due to the confinement imposed by the bubble trap. Another aspect of our study explores the topological change from a bubble trap to a filled BEC, demonstrating that tuning the trap parameter $\Delta$ can serve as a protocol to generate higher angular momentum states in harmonic confinement. This procedure has the potential to bring the interacting atomic gas into the desired quantum-Hall regime [92, 93]. The persistent current and hypersonic velocity exhibited by atomic dipolar BECs in a rotating bubble trap could pave the way for their utilization in quantum transport applications within atomtronics [94, 95, 96].
Acknowledgment.- We thank Vito Scarola for carefully reading the manuscript and insightful discussions. We also thank Arpana Saboo for the helpful discussions. We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of PARAM Shakti at IIT Kharagpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. H. S. Ghosh gratefully acknowledges the support from the Prime Minister’s Research Fellowship (PMRF), India. S. Halder acknowledges the MHRD Govt. of India for the research fellowship. S. Das acknowledges support from AFOSR FA9550-23-1-0034.